# Two variants of the friendship paradox: The condition for inequality between them

### v2

@@ I Introduction
The friendship paradox (FP), first articulated by Feld [1], states that the average degree of an individual’s friends typically exceeds the individual’s own degree. This phenomenon arises because high-degree nodes are disproportionately represented when sampling along edges. In practice, two closely related but distinct average quantities are used. The alter-based mean computes the expected degree encountered by following a uniformly random edge. The ego-based mean first computes, for each node, the average degree of its neighbors and then averages that quantity over nodes. Both are natural formalizations of “how popular one’s friends are on average,” yet they need not coincide outside special cases such as regular or degree-uncorrelated networks.
@@ II Tale of Two FP Types
    This quantity is greater than or equal to $\langle k\rangle$, yielding the classical form of the FP: $\langle k_{\text{friend}}\rangle_{\mathrm{n}}\geq\langle k\rangle_{\mathrm{n}}$.
@@
In the terminology of Ref. [2], $\langle k_{\text{friend}}\rangle_{\mathrm{n}}$ corresponds to the alter-based mean and $\langle k_{nn}\rangle_{\mathrm{n}}$ to the ego-based mean. Although they coincide in regular or degree-uncorrelated networks, their difference in general encodes the structural correlation between a node’s degree and that of its neighbors [3]. The remainder of this document analyzes this difference and establishes its precise relationship to the degree-degree covariance for each node and its nearest neighbors.
@@ IV Examples
Figure 1: A simple five-node network where the alter-based and ego-based means coincide (neutral case). The network values are: $\langle k\rangle_{\mathrm{n}}=2$, $\langle k_{\text{friend}}\rangle_{\mathrm{n}}=11/5$, $\langle k_{nn}\rangle_{\mathrm{n}}=11/5$, and $\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})=0$.
This small example represents a neutral structure: higher- and lower-degree nodes coexist without systematic degree–degree correlation. As a result, the FP appears identically under both the alter-based and ego-based averages, illustrating $\langle k_{\text{friend}}\rangle_{\mathrm{n}}=\langle k_{nn}\rangle_{\mathrm{n}}$.
@@
Figure 2: A network with mild assortativity: high-degree nodes tend to connect to one another. The network values are: $\langle k\rangle_{\mathrm{n}}=7/3$, $\langle k_{\text{friend}}\rangle_{\mathrm{n}}=18/7$, $\langle k_{nn}\rangle_{\mathrm{n}}=5/2$, and $\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})=1/6$.
$\langle k_{\text{friend}}\rangle_{\mathrm{n}}-\langle k_{nn}\rangle_{\mathrm{n}}=1/14=\frac{1}{\langle k\rangle_{\mathrm{n}}}\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})\,.$
Here the positive covariance indicates assortative mixing, in which high-degree nodes preferentially attach to other high-degree nodes. Consequently, the edge-weighted (alter-based) average $\langle k_{\text{friend}}\rangle_{\mathrm{n}}$ exceeds the node-weighted (ego-based) average $\langle k_{nn}\rangle_{\mathrm{n}}$, consistent with Eq. (13).
@@
Figure 3: A star-like configuration exhibiting strong disassortativity. The network values are: $\langle k\rangle_{\mathrm{n}}=8/5$, $\langle k_{\text{friend}}\rangle_{\mathrm{n}}=5/2$, $\langle k_{nn}\rangle_{\mathrm{n}}=17/5$, and $\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})=-36/25$.
$\langle k_{\text{friend}}\rangle_{\mathrm{n}}-\langle k_{nn}\rangle_{\mathrm{n}}=-9/10=\frac{1}{\langle k\rangle}\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})\,.$
The central node’s high degree paired with peripheral nodes of degree one creates a clear disassortative structure. Because low-degree nodes dominate the population, the node-averaged mean of neighbor degrees ($\langle k_{nn}\rangle_{\mathrm{n}}$) becomes larger than the edge-averaged mean, yielding a negative covariance, the opposite limit of Fig. 2. A simple example is the complete bipartite network composed of $n_{1}$ and $n_{2}$ nodes in the bipartite groups and all of the $n_{1}n_{2}$ possible connections between the two groups exist, respectively, where
@@ VI Discussion and conclusions
We have demonstrated that the covariance form presented in this document and the moment–based expression of Ref. [2] describe the same underlying structural relation, albeit from two complementary statistical viewpoints. The inversity parameter $\rho$ is defined as the Pearson correlation on the oriented-edge ensemble between a node’s degree and the inverse degree of its neighbor:

### v3

@@ I Introduction
The friendship paradox (FP), first articulated by Feld [1], states that the average degree of an individual’s friends typically exceeds the individual’s own degree. This phenomenon arises because high-degree nodes are disproportionately represented when sampling along edges. In practice, two closely related but distinct average quantities are used. The alter-based mean computes the expected degree encountered by following a uniformly random edge. The ego-based mean first computes, for each node, the average degree of its neighbors and then averages that quantity over nodes. Both are natural formalizations of “how popular one’s friends are on average,” yet they need not coincide outside special cases such as regular or degree-uncorrelated networks. Beyond these average-based statements, recent work has further clarified how majority-type versions of the paradox—often phrased as ‘most people’s friends have more friends’—can be different from the classical FP. In a recent paper [2], I analyze this so-called strong friendship paradox [3, 4, 5, 6] and show how local degree structure can reverse or amplify the classical expectation.
@@ II Tale of Two FP Types
    This quantity is greater than or equal to $\langle k\rangle_{\mathrm{n}}$, yielding the classical form of the FP: $\langle k_{\text{friend}}\rangle_{\mathrm{n}}\geq\langle k\rangle_{\mathrm{n}}$.
@@
In the terminology of Ref. [7], $\langle k_{\text{friend}}\rangle_{\mathrm{n}}$ corresponds to the alter-based mean and $\langle k_{nn}\rangle_{\mathrm{n}}$ to the ego-based mean. Although they coincide in regular or degree-uncorrelated networks, their difference in general encodes the structural correlation between a node’s degree and that of its neighbors [8]. The remainder of this paper analyzes this difference and establishes its precise relationship to the degree-degree covariance for each node and its nearest neighbors.
@@ IV Examples
(a) (b) (c)
Figure 1: (a) A simple five-node network where the alter-based and ego-based means coincide (neutral case). The network values are: $\langle k\rangle_{\mathrm{n}}=2$, $\langle k_{\text{friend}}\rangle_{\mathrm{n}}=11/5$, $\langle k_{nn}\rangle_{\mathrm{n}}=11/5$, and $\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})=0$. (b) A network with mild assortativity: high-degree nodes tend to connect to one another. The network values are: $\langle k\rangle_{\mathrm{n}}=7/3$, $\langle k_{\text{friend}}\rangle_{\mathrm{n}}=18/7$, $\langle k_{nn}\rangle_{\mathrm{n}}=5/2$, and $\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})=1/6$. (c) A star-like configuration exhibiting strong disassortativity. The network values are: $\langle k\rangle_{\mathrm{n}}=8/5$, $\langle k_{\text{friend}}\rangle_{\mathrm{n}}=5/2$, $\langle k_{nn}\rangle_{\mathrm{n}}=17/5$, and $\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})=-36/25$.
The example in Fig. 1(a) represents a neutral structure: higher- and lower-degree nodes coexist without systematic degree–degree correlation. As a result, the FP appears identically under both the alter-based and ego-based averages, illustrating $\langle k_{\text{friend}}\rangle_{\mathrm{n}}=\langle k_{nn}\rangle_{\mathrm{n}}$.
@@
In the example in Fig. 1(b), the positive covariance indicates assortative mixing, in which high-degree nodes preferentially attach to other high-degree nodes. Consequently, the edge-weighted (alter-based) average $\langle k_{\text{friend}}\rangle_{\mathrm{n}}$ exceeds the node-weighted (ego-based) average $\langle k_{nn}\rangle_{\mathrm{n}}$, consistent with Eq. (13):
$\langle k_{\text{friend}}\rangle_{\mathrm{n}}-\langle k_{nn}\rangle_{\mathrm{n}}=\frac{1}{14}=\frac{1}{\langle k\rangle_{\mathrm{n}}}\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})\,.$
@@
In the example in Fig. 1(c), the central node’s high degree paired with peripheral nodes of degree one creates a clear disassortative structure. Because low-degree nodes dominate the population, the node-averaged mean of neighbor degrees ($\langle k_{nn}\rangle_{\mathrm{n}}$) becomes larger than the edge-averaged mean, yielding a negative covariance, the opposite limit of Fig. 1(b):
$\langle k_{\text{friend}}\rangle_{\mathrm{n}}-\langle k_{nn}\rangle_{\mathrm{n}}=-\frac{9}{10}=\frac{1}{\langle k\rangle}\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})\,.$
A simple example is the complete bipartite network composed of $n_{1}$ and $n_{2}$ nodes in the bipartite groups and all of the $n_{1}n_{2}$ possible connections between the two groups exist, respectively, where
@@ VI Discussion and conclusions
We have demonstrated that the covariance form presented in this paper and the moment–based expression of Ref. [7] describe the same underlying structural relation, albeit from two complementary statistical viewpoints. The inversity parameter $\rho$ is defined as the Pearson correlation on the oriented-edge ensemble between a node’s degree and the inverse degree of its neighbor: